OpenStudy (anonymous):

Joey's sock drawer is unorganized and contains 3 black dress socks, 5 black ankle socks, 7 brown dress socks, and 6 brown ankle socks. What is the probability that Joey will blindly reach into his sock drawer and pull out a sock that is brown or a dress sock?

OpenStudy (anonymous):

total num of items = 3+5+7+6=23 so sample space = 23. Among them 13 are brown and 10 are dress socks. The question says "or" so we should first find the probability of having a brown sock (which is 13/23), add the probability having a dress sock to it (which is 10/23) and subtract the intersection of two (which is brown dress = 7/23) so the answer would be 16/23

OpenStudy (anonymous):

All of the items added up is 21 though. @barbillus so the answer is actually 16/21, correct?

OpenStudy (anonymous):

yeah sorry, that's 21, my mistake